### v1

# Efficient Tactile Perception with Soft Electrical Impedance Tomography and Pre-trained Transformer

Abstract
Tactile sensing is fundamental to robotic systems, enabling interactions through physical contact in multiple tasks. Despite its importance, achieving high-resolution, large-area tactile sensing remains challenging. Electrical Impedance Tomography (EIT) has emerged as a promising approach for large-area, distributed tactile sensing with minimal electrode requirements which can lend itself to addressing complex contact problems in robotics. However, existing EIT-based tactile reconstruction methods often suffer from high computational costs or depend on extensive annotated simulation datasets, hindering its viability in real-world settings. To address this shortcoming, here we propose a Pre-trained Transformer for EIT-based Tactile Reconstruction (PTET), a learning-based framework that bridges the simulation-to-reality gap by leveraging self-supervised pretraining on simulation data and fine-tuning with limited real-world data. In simulations, PTET requires 99.44% fewer annotated samples than equivalent state-of-the-art approaches (2,500 vs. 450,000 samples) while achieving reconstruction performance improvements of up to 43.57% under identical data conditions. Fine-tuning with real-world data further enables PTET to overcome discrepancies between simulated and experimental datasets, achieving superior reconstruction and detail recovery in practical scenarios. PTET’s improved reconstruction accuracy, data efficiency, and robustness in real-world tasks establish it as a scalable and practical solution for tactile sensing systems in robotics, especially for object handling and adaptive grasping under varying pressure conditions.
Index Terms:
Masked Autoencoder, Electrical Impedance Tomography, Tactile Sensing, Human-Machine Interaction.
I Introduction
Tactile sensing is a foundational technology in robotics, enabling robots to effectively interact with their environment by perceiving pressure, texture, and force through physical contact [1]. This capability is essential for tasks such as object manipulation, human-robot interaction, and environment exploration [2, 3, 4]. Despite its significance, achieving high-resolution, large-area tactile sensing remains a significant challenge due to inherent trade-offs among accuracy, scalability, and complexity in sensor design. Current tactile sensors typically rely on capacitive [5], piezoresistive [6], or piezoelectric [7, 8] mechanisms. However, the spatial resolution of these systems is constrained by the number of sensor array pixels, which require densely packed sensor arrays for higher spatial resolution. This not only complicates fabrication and increases costs but also limits flexibility and scalability for expanded tactile sensing tasks [9].
To overcome these limitations, Electrical Impedance Tomography (EIT) has emerged as a promising alternative [10]. EIT offers several advantages over traditional methods, including the ability to achieve large-area tactile sensing with fewer boundary electrodes [11, 12, 13]. By injecting small electrical currents through electrode pairs and measuring the resulting voltage changes, EIT reconstructs the conductivity distribution of a conductive soft layer to infer tactile inputs such as pressure or deformation [14, 15]. However, EIT tactile reconstruction poses significant challenges in terms of accuracy and computational efficiency [16].
Model-based methods [17, 18, 19, 20] have been typically employed for EIT reconstruction, solving the inverse problem non-iteratively or iteratively to estimate the conductivity distribution from voltage measurements. While these approaches do not require pre-collected datasets, they often produce suboptimal reconstruction results, especially for dynamic tactile sensing tasks. This limitation is largely attributed to EIT’s inherently ill-posed and ill-conditioned inverse problem [16], exacerbating reconstruction quality issues. Furthermore, the accuracy of model-based approaches is inherently limited due to simplified assumptions on the governing physical model, sensor geometry and materials properties [21].
In recent years, data-driven approaches, particularly those based on deep learning, have gained traction in the EIT domain [22, 23, 24]. These methods aim to train deep nets to directly map voltage measurements to tactile reconstructions, bypassing the iterative processes and linear approximations. While data-driven approaches have shown promise in delivering faster and more accurate reconstructions, they are heavily dependent on large, annotated datasets [25, 26, 27, 28]. Acquiring extensive real-world datasets with ground-truth tactile annotations ($e.g.,>100k$) is often impractical due to the labour-intensive and time-consuming nature of data collection. Consequently, these methods frequently depend on simulation-generated data for model training. However, the inherent discrepancies between simulation and experimental data can introduce considerable reconstruction errors in practical applications. These discrepancies manifest themselves as reconstruction errors, including inaccuracies in pressure magnitude, the appearance of spurious residual artefacts, and misrepresentation of the shape of the contact interface in complex scenarios [29, 30].
To address this issue, self-supervised pretraining has emerged as a powerful technique, enabling models to extract general features from unlabelled data and subsequently transfer the features to various downstream tasks by supervised fine-tuning using limited labelled data. This paradigm has demonstrated success in robotics applications, such as control [31], navigation [32], visual perception [33], and sim-to-real transfer [34]. Recent advancements in self-supervised learning have particularly focused on masked prediction in large models like GPT [35], and BERT [36], showcasing their potential in stacking large language/vision models. For high-dimensional data, one pivotal work in computer vision is the masked autoencoder (MAE) [37], which conducts a scalable vision learner to extract image features. Inspired by this, EIT can similarly be pre-trained in an unlabelled manner using a self-supervised approach. This enables EIT tactile reconstruction models to learn from unlabelled data and subsequently fine-tuned with limited paired (labelled) real-world data, establishing a real-to-real learning pipeline that eliminates the need for extensive real-world datasets while mitigating simulation-to-reality discrepancies.
Figure 1: Electrical Impedance Map (EIM). (a) EIT-based tactile sensing principle. (b) Original EIT measurements. (c) Construction of EIM: an example. (d) Enhanced EIM (E2IM).
In this paper, we propose the Pre-trained Transformer for the EIT-based Tactile Reconstruction (PTET) model, a novel framework that addresses two key challenges in learning-based EIT tactile sensing: (1) the heavy dependency on large annotated datasets and (2) the simulation-to-reality gaps that hinder practical applications. PTET leverages a self-supervised pretraining strategy to learn generalizable features from unlabelled tactile sensing data, significantly reducing reliance on large annotated datasets. The model is subsequently fine-tuned using a limited amount of annotated real-world data, enabling more accurate tactile reconstruction while mitigating inaccuracies typically introduced by simulation-based training. This dual-phase learning strategy allows PTET to bridge the gap between experimental conditions and practical applications. Unlike traditional learning-based approaches that require extensive labelled datasets, which are typically generated from simulation, PTET effectively operates with a small fraction of the real-world data, making it scalable and practical for real-world deployments. The main contributions of this work are as follows:
  • We introduce the PTET model, which leverages large-scale self-supervised pretraining combined with few-shot fine-tuning to reduce dependency on large annotated datasets. This approach achieves more accurate tactile reconstructions with significantly fewer samples compared to traditional learning-based methods, making it more suitable for practical tactile systems.
  • We develop a flexible, large-area EIT-based tactile sensor that introduces a dual-conductivity layer structure, uniquely combining hydrogel for flexibility and biocompatibility with a carbon black-to-graphite (CBG) composite for significantly enhanced conductivity and mechanical flexibility. Compared to existing designs, our sensor achieves superior signal stability and high sensitivity, capable of detecting compressive strain as small as 1 mm over a large sensing area.
  • For the first time, we demonstrate high-resolution tactile reconstruction using a limited experimental dataset in a learning-based framework. This real-to-real approach eliminates errors and mismatches associated with simulated data, making EIT-based tactile sensing more practical and robust for real-world applications.
II Enhanced Electrical Impedance Map
In each EIT measurement, two neighbouring electrodes are used for current injection, while the remaining electrodes are selected for voltage measurements. Among the measuring electrodes, measurements are taken between neighbouring electrode pairs, excluding the current injection electrodes (see Fig. 1a). For the EIT-based tactile sensor with 16 electrodes, this results in 16 (excitation electrode pairs) × 13 (measurement electrode pairs) = 208 measurements per frame, though only 104 independent voltages are recorded due to the reciprocity theorem.
To capture the spatial and geometric features of the EIT measurements in an image-like format [38], we convert the original voltage measurements (104 $\times$1 vector, Fig. 1b) to a 16 $\times$ 16 matrix, termed the Electrical Impedance Map (EIM) (Fig. 1c). The EIM matrix is constructed in Algorithm 1.
Algorithm 1 Construction of EIM
1:$\mathbf{v}_{\text{ori}}$: Original voltage measurements (104 $\times$ 1 vector)
2:$\mathbf{v}_{\text{eim}}$: EIM ($16\times 16$ matrix)
3:Initialize $\mathbf{v}_{\text{eim}}\leftarrow\mathbf{0}_{16\times 16}$triangleright Zero matrix of size $16\times 16$
4:$index\leftarrow 1$
5:for $i=1$ to $16$ do
6:     for $j=1$ to $16$ do
7:         if $j>i+1$ and $index\leq 104$ then
8:              $\mathbf{v}_{\text{eim}}(i,j)\leftarrow\mathbf{v}_{\text{ori}}[index]$
9:              $\mathbf{v}_{\text{eim}}(j,i)\leftarrow\mathbf{v}_{\text{ori}}[index]$triangleright Apply reciprocity theorem
10:              $index\leftarrow index+1$
11:         end if
12:     end for
13:end for
14:return $\mathbf{v}_{\text{eim}}$
To further enhance the representation quality of EIM, we propose using Bicubic interpolation to enhance the EIM to the Enhanced-EIM (E2IM) (following the Algorithm 2). Unlike the original $16\times 16$ matrix where each voltage measurement occupies a single pixel, the E2IM represents each measurement as a $4\times 4$ patch, generating a $64\times 64$ matrix (Fig. 1d). This augmentation improves the spatial granularity and feature richness of the original EIM, enhancing the model’s capability to capture fine-grained spatial details and structural patterns, essential for achieving high-resolution tactile reconstructions and improving the robustness of the learning process.
Algorithm 2 Construction of E2IM
1:$\mathbf{v}_{\text{eim}}$: EIM ($16\times 16$ matrix)
2:$\mathbf{v}$: E2IM ($64\times 64$ matrix)
3:Initialize $\mathbf{v}\leftarrow\mathbf{0}_{64\times 64}$triangleright Zero matrix of size $64\times 64$
4:Apply bicubic interpolation to $\mathbf{v}_{\text{eim}}$ to resize it from $16\times 16$ to $64\times 64$
5:Assign the resized matrix to $\mathbf{v}$
6:return $\mathbf{v}$
III PTET: Pre-trained Transformer for EIT-based Tactile Reconstruction
Figure 2: Overview of PTET. Our method comprises three main processes: bottom— 1) Pretraining with E2IM, for an effective and robust voltage encoder $\mathcal{E}_{vol}$, through self-supervised learning by predicting masked regions in grey. 2) Pretraining with tactile map, through self-supervised learning, for a robust tactile decoder $\mathcal{D}_{tact}$; top— 3) Fine-tuning with limited pairs from real-world measurements, the fine-tuned model will be deployed for reconstruction. Notably, during fine-tuning, $\mathcal{E}_{vol}$’s parameters are frozen, while $\mathcal{D}_{tact}$’s are tunable.
The reconstruction of tactile representations t from EIT voltage measurements v, represented in the format of E2IM, can be described by a direct mapping $f$:
$f:\textbf{v}\rightarrow\textbf{t}.$ (1)
To enhance this transformation, we introduce a latent variable z, enabling the process to be decomposed into two sequential mappings:
$f:\textbf{v}\rightarrow\textbf{z}\rightarrow\textbf{t},$ (2)
where $p_{v}:\textbf{v}\rightarrow\textbf{z}$ maps the voltage data to the latent variable, and $p_{t}:\textbf{z}\rightarrow\textbf{t}$ maps the latent variable to tactile representations. This decomposition compresses the information from the voltage and tactile spaces into a lower-dimensional representation, called latent space, allowing for effective feature extraction during the independent learning of $p_{v}$ and $p_{t}$. By leveraging self-supervised pretraining, this approach allows the models $p_{v}$ and $p_{t}$ to independently learn representations from the unlabeled data v and t, respectively, with subsequent fine-tuning focused on aligning the latent variables.
To implement this approach, we propose the PTET framework, built on a pretraining-based auto-encoding architecture, as illustrated in Fig.2. PTET comprises three key components:
  1. 1.
    Pretraining with E2IM: In this stage, a masked autoencoder $\mathcal{D}_{vol}(\mathcal{E}_{vol}(\textbf{v}))$ is employed to extract features from unlabeled voltage data v, which is represented by E2IM. This process trains an effective and robust encoder $\mathcal{E}_{vol}(\textbf{v})$, functioning mapping $p_{v}:\textbf{v}\rightarrow\textbf{z}_{v}$, which compresses the input v to a latent variable $\textbf{z}_{v}$;
  2. 2.
    Pretraining with tactile map: In parallel with 1), an image autoencoder $\mathcal{D}_{tact}(\mathcal{E}_{tact}(\textbf{t}))$ is utilized to learn perceptual features from unlabeled tactile data t, represented by the 2D tactile map. This process trains a robust tactile map decoder $\mathcal{D}_{tact}(\textbf{t})$, generating the target output t from a latent variable $\textbf{z}_{t}$ to implement the mapping $p_{t}$;
  3. 3.
    Fine-tuning with limited paired voltage-tactile data: Utilizing the pretrained models from the above two steps, the voltage encoder $\mathcal{E}_{vol}$ and tactile decoder $\mathcal{D}_{tact}$ are aligned in a fine-tuned model, $\mathcal{D}_{tact}(\mathcal{E}_{vol}(\textbf{v}))$, through supervised training using limited paired v-t data. This alignment focuses on the mapping between latent variables: $\textbf{z}_{v}\rightarrow\textbf{z}_{t}$, enabling the final model to reconstruct tactile maps from the E2IM.
The PTET framework adopts a dual-phase training scheme: the pretraining phase with 1) and 2) and fine-tuning phase with 3). In the pretraining phase, the functioning $\mathcal{E}_{vol}$ and $\mathcal{D}_{tact}$ are trained with large-scale unannotated datasets. It allows the model to learn generalizable features from E2IM and tactile maps separately based on a large scale of random simulation data, which can be readily acquired. While in the fine-tuning phase, using only limited paired data from real-world measurements, the pretrained models are aligned to establish the relationship between the voltage and tactile domains. The tactile map can then be reconstructed from a given E2IM describing real-world voltage measurements. Notably, in both 1) and 2), PTET employs an asymmetric architecture for encoders and decoders to enhance the efficiency and effectiveness of feature representation in $\mathcal{E}_{vol}$ and $\mathcal{D}_{tact}$, as detailed in the following sections.
III-A Pretraining with E2IM
III-A1 Overview
To pretrain E2IM v, we designed a Transformer-based autoencoder, consisting of voltage encoder $\mathcal{E}_{vol}$ and tactile decoder $\mathcal{D}_{tact}$. As shown in Fig. 2, the voltage autoencoder trains the input E2IM v in a self-supervised loop $\textbf{v}\rightarrow\hat{\textbf{v}}$. It first maps the input v to the latent representation $\textbf{z}_{v}$:
$\textbf{z}_{v}=\mathcal{E}_{vol}(\textbf{v}),$ (3)
then, the decoder maps the latent representation $\textbf{z}_{v}$ back to the reconstructed E2IM, $\hat{\textbf{v}}$:
$\hat{\textbf{v}}=\mathcal{D}_{vol}(\textbf{z}_{v}).$ (4)
The pretraining loop $\textbf{v}\rightarrow\textbf{z}_{v}\rightarrow\hat{\textbf{v}}$ can be framed as an optimization problem that minimizes the difference between the original E2IM v and the reconstructed output $\hat{\textbf{v}}$. The loss function $\mathcal{L}_{vol}$ is computed as the mean squared error (MSE) between the input and the reconstructed E2IM:
$\arg\min_{\mathcal{E}_{vol},\mathcal{D}_{vol}}\mathcal{L}_{vol}=\|\hat{\textbf{v}}-\textbf{v}\|^{2}.$ (5)
This objective function aims to minimize the reconstruction error by optimizing the voltage encoder $\mathcal{E}_{vol}$ and decoder $\mathcal{D}_{vol}$.
III-A2 Masked autoencoder framework
The pretraining model for E2IM utilizes the masked autoencoder framework [37] based on vision Transformer (ViT) [39]. The key element for this model is the Transformer layer [40], which processes data in sequential tokens. As described in Section II, voltage data v is represented by E2IM, a $64\times 64$ matrix. To process this form, E2IM data is pre-processed through masking and tokenization, as illustrated in Tokenization in Fig. 2. Specifically, for a batch of input E2IM matrices, $\textbf{v}\in\mathbb{R}^{64\times 64}$ is reshaped into a sequence of flattened 2D patches $\textbf{v}_{p}\in\mathbb{R}^{N\times(P^{2})}$. Here, $(P,P)=4^{2}$ is the resolution of each patch, $N=\frac{64\times 64}{4^{2}}=16\times 16$ is the resulting number of the patches.
Then, these patches $\textbf{v}_{p}$, preceded by an instructive class token $\textbf{v}_{cls}$ at the start of the sequence as $\textbf{v}_{p}^{0}$, are linearly projected into the Transformer layer and combined with patch and positional embeddings [39] for tokenization. The resulting input sequence, denoted as voltage token $\textbf{z}_{0}$, is computed as follows:
$\textbf{z}_{0}=[\textbf{v}_{cls};\textbf{v}_{p}^{1}\textbf{E};\textbf{v}_{p}^{2}\textbf{E};\cdots;\textbf{v}_{p}^{N}\textbf{E}]+\textbf{E}_{pos},$ (6)
where $\textbf{E}\in\mathbb{R}^{N}$ is the patch embedding vector, $N=16\times 16$ is the total number of patches, and $\textbf{E}_{pos}\in\mathbb{R}^{N+1}$ represents the positional embedding vector.
The mask sampling strategy is applied to train the voltage encoder $\mathcal{E}_{vol}$. Specifically, a high masking ratio ($\alpha=75\%$ [37]) randomly masks $4\times 4$ $\textbf{v}_{p}\textbf{E}$ tokens globally. This approach maximizes information entropy and sets the masked tokens as prediction targets, enhancing the model’s learning generalizability. It leaves only the unmasked voltage tokens $(1-\alpha)\textbf{z}_{0}$ as input to the Transformer encoder $\mathcal{E}_{vol}$ to extract features:
$\textbf{z}_{1}=[\textbf{v}_{cls};\textbf{v}_{p}^{1}\textbf{E};\textbf{v}_{p}^{2}\textbf{E};\cdots;\textbf{v}_{p}^{(1-\alpha)N}\textbf{E}]+\textbf{E}_{pos}.$ (7)
Transformer layers process the $\textbf{z}_{1}$ with layers of multi-head attention [40] ($\mathcal{A}$) with Layer Normalization ($LN$):
$\textbf{z}_{l}^{\prime}=\mathcal{A}(LN(\textbf{z}_{l-1}))+\textbf{z}_{l-1},$ (8)
where $l=2\cdots(L+1)$ denotes the $l_{th}$ layer, $L=(1-\alpha)N$ represents the total number of unmasked voltage tokens $\textbf{v}_{p}$. Then the outputs are processed by multi-layer perception (MLP) as follows:
$\textbf{z}_{l}=MLP(LN(\textbf{z}_{l}^{\prime}))+\textbf{z}_{l}^{\prime},$ (9)
and the encoded latent output $\textbf{z}_{v}$ from $\mathcal{E}_{vol}$ is:
$\textbf{z}_{v}=LN(\textbf{z}_{l}).$ (10)
To complete Equation 4, the voltage decoder $\mathcal{D}_{vol}(\textbf{z}_{v})$ utilizes positional embedding information $\textbf{E}_{pos}$ retained during tokenization to predict masked tokens and re-position them in the E2IM image, as shown in Fig. 2. The predicted masked region combines the unmasked part with the reconstructed E2IM $\hat{\textbf{v}}$. Unlike previous methods that processed voltage data v sequentially [27] and ignored the spatial arrangement of sensors, our approach employs a 2D E2IM matrix representation. This 2D form enables the PTET model to capture and save spatial information associated with electrodes in $\textbf{E}_{pos}$, ensuring the positional information is maintained throughout the learning process.
III-A3 Implementation
We designed the voltage encoder $\mathcal{E}_{vol}$ and decoder $\mathcal{D}_{vol}$ with customized scales and hyperparameters to match the characteristics of the E2IM input. To effectively capture information, the embedding dimension was optimized and set to 256, aligning with the resolution of the E2IM in $64\times 64$. The architecture adopts an asymmetric design to enhance the encoder’s feature extraction capacity while simplifying the decoder. Specifically, the voltage encoder $\mathcal{E}_{vol}$ consists of 12 Transformer layers, whereas the voltage decoder $\mathcal{D}_{vol}$ is deliberately kept shallower with only 2 Transformer layers. All Transformer layers are configured with 4 attention heads to maintain architectural consistency.
III-A4 E2IM’s enhancement on pretraining
The study of MAE [37] has shown that mask sampling strategies significantly impact model performance, although the underlying reasons were not explored in depth. In MAE, an ablation study was conducted on highly semantic images from ImageNet[41], where the less semantic structural information masked, the easier the image could be reconstructed.
However, as depicted in Fig. 1, EIM represents a low-semantic representation of the voltage. In this context, applying large patch masking obscures key structural information, leading to poor reconstruction. Conversely, applying small patch masking reduces the representation quality, which hampers effective feature extraction. For example, $1\times 1$ masking will allow reconstruction through interpolation but limits the model’s ability to learn meaningful representations. Therefore, the minimum and optimal patch size $P$ for a $16\times 16$ EIM is $2$.
As we introduced before, to address this issue, we extended EIM from $16\times 16$ to $64\times 64$ and set the patch resolution to $4\times 4$, enhancing the model’s ability to learn and represent features effectively. To validate the effectiveness of the E2IM, we conducted ablation studies comparing its performance against the original EIM in terms of reconstructions in pertaining. As summarized in Table. I, MSE and Structural Similarity Index Measure (SSIM) were calculated using the mean values over 25,000 samples. The results demonstrate that E2IM achieves a 99.95% improvement in MSE and a 66.07% increase in SSIM over EIM, underscoring its superior representation quality.
TABLE I: Encoder pretraining with EIM and E2IM
EIM E2IM
Dimension 16×16 64×64
$P$ 2 4
MSE 5.6466e-4 2.9701e-7
SSIM 0.5991 0.9949
III-B Pretraining with tactile maps
III-B1 Framework
The tactile map t is set as $48\times 48$, i.e., $\textbf{t}\in\mathbb{R}^{48\times 48}$. Similar to $\mathcal{D}_{vol}(\mathcal{E}_{vol}(\textbf{v}))$, we designed a tactile autoencoder based on ResNet [42] for self-supervised pretraining. The tactile autoencoder model consists of an encoder $\mathcal{E}_{tact}$ and a decoder $\mathcal{D}_{tact}$. The pretraining of tactile map $\textbf{t}\rightarrow\textbf{z}_{t}\rightarrow\hat{\textbf{t}}$ can be presented as:
$\arg\min_{\mathcal{E}_{tact},\mathcal{D}_{tact}}\mathcal{L}_{tact}\|\hat{\textbf{t}}-\textbf{t}\|^{2},$ (11)
where $\mathcal{L}_{tact}$ represents the MSE between the input and the reconstructed tactile map. The encoder $\mathcal{E}_{tact}$ encodes t into a latent representation $\textbf{z}_{t}$:
$\textbf{z}_{t}=\mathcal{E}_{tact}(\textbf{t}),$ (12)
The decoder $\mathcal{D}_{tact}$ predicts from $\textbf{z}_{t}$ to finish the self-supervised learning cycle:
$\hat{\textbf{t}}=\mathcal{D}_{tact}(\textbf{z}_{t}).$ (13)
III-B2 Implementation
The tactile autoencoding model $\mathcal{D}_{tact}(\mathcal{E}_{tact}(\textbf{t}))$ is stacked with convolutional ResNet blocks. Each convolutional ResNet block consists of two ResNet layers followed by a convolution layer. These blocks perform downsampling or upsampling, with the number of channels in each layer ranging from [32, 64, 128].
Similar to the asymmetric strategy in $\mathcal{E}_{vol}$, we integrated the residual self-attention layer [40, 43] only into the decoder $\mathcal{D}_{tact}$ to enhance the tactile reconstruction capacity from the latent variable $\textbf{z}_{t}$. In $\mathcal{D}_{tact}$, a residual self-attention layer is integrated between the first two ResNet layers to enhance feature representation. The self-attention mechanism is implemented using three separate 128-dimensional linear layers to compute the Query ($\mathbf{Q}$), Key ($\mathbf{K}$), and Value ($\mathbf{V}$) matrices [40]. The attention weights are calculated as follows:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (14)
where $d_{k}=128$ is the dimensionality of the Query and Key vectors. This layer refines feature extraction by selectively focusing on relevant spatial and semantic information while preserving residual connections for stable training.
III-C Fine-tuning with limited volatge-tactile pairs
One of the primary challenges in learning-based tactile sensing lies in the scarcity of annotated real-world data for training. To address this limitation, we employ fine-tuning by integrating the pretrained voltage encoder $\mathcal{E}_{vol}$ and the pretrained tactile decoder $\mathcal{D}_{tact}$ into a unified model. This model is then refined using a limited dataset of voltage-tactile paired samples, enabling effective adaptation to the target task despite the constrained data availability. Due to the vast differences in scale between the small fine-tuning datasets (i.e., down to 2500 paired training sets) and large pretraining datasets (i.e., 450,000 unlabeled training sets), it often leads to over-fitting. To solve this problem, we propose strategies for few-shot fine-tuning for tiny datasets as follows:
III-C1 Latent space compression
As outlined in Sections III-A and III-B, the deep neural network architecture of the autoencoder achieves a high compression ratio by encoding voltage and tactile data into a low-dimensional latent space. This compressed representation facilitates efficient data transformation and adaptation during the fine-tuning process.
III-C2 Partial tuning
we observed significant differences in the number of parameters between the voltage encoder $9.54M$ and the tactile decoder $1.23M$. Therefore, as shown in Fig. 2, we freeze all parameters in the voltage encoder $\mathcal{E}_{vol}$ and only update the parameters in the tactile decoder $\mathcal{D}_{tact}$ during the fine-tuning stage, as shown in Fig.2. The latent variables from voltage measurement conditioning $\textbf{z}_{v}$ and tactile $\textbf{z}_{t}$ are connected via direct linear projection without any hidden layers:
$\textbf{z}_{t}=MLP(\textbf{z}_{v});$ (15)
$\mathcal{D}_{tact}$ is therefore updated by:
$\arg\min_{\mathcal{D}_{tact}}\mathcal{L}_{tact}\|\mathcal{D}_{tact}(\textbf{z}_{t})-\textbf{t}\|^{2},$ (16)
where the loss function $\mathcal{L}_{tact}$, the same as in tactile map pretraining, is calculated by MSE between the reconstructed tactile map $\hat{\textbf{t}}=\mathcal{D}_{tact}(\textbf{z}_{t})$ and the ground truth tactile map t.
IV PTET performance on simulated test
IV-A Dataset
Figure 3: Example of a simulation sample. (a) E2IM (b) Ground Truth (GT) of the tactile pattern. (c) GT after Gaussian filtering. Note that we use (b) in model training and quantitative evaluation and (c) for visualization.
As shown in Fig. 3, the sensing region, measuring $100\times 100\,mm^{2}$, is divided into $48\times 48$ pixels with a background conductivity of 0.00312 S/m. Circular areas with diameters ranging from 7.5 mm to 27.5 mm are used to represent tactile interactions, emphasizing both position and intensity. The conductivity within these touch regions is randomly varied between 0.05 and 2 times the background value to simulate diverse touch effects. The dataset generated comprises 500,000 samples, with each of the five subsets containing 100,000 samples. Each sample includes one E2IM and corresponding Ground Truth (GT). These subsets correspond to different numbers of touch regions, ranging from 1 to 5 touch areas per sample. Each subset is further partitioned into training, validation, and testing datasets using an 18:1:1 ratio. This results in a total of 450,000 samples for training, 25,000 for validation, and 25,000 for testing. All data were generated using COMSOL Multiphysics and Matlab.
IV-B Comparison Algorithms
To evaluate the effectiveness of PTET, we compared it with the state-of-the-art (SOTA) DNN-based algorithm [28], another autoencoder-based model designed for EIT-based tactile sensing. Additionally, we developed a supervised variant of PTET, termed PTET-SL, which employs the same neural network architecture as the fine-tuned model but is trained directly on paired datasets without pretraining. This modification aims to demonstrate the reliance of supervised models on large datasets while showcasing the superior performance of PTET. For a fair comparison, all algorithms utilized a unified mean squared error (MSE) loss function. The parameters for the SOTA algorithm were the same as in [28], with only minor adjustments made to accommodate the input dimensions of our dataset. Both PTET and PTET-SL were trained using identical parameters (see Table II).
TABLE II: Training parameters
Parameter Pretraining for $\mathcal{E}_{vol}$ Pretraining for $\mathcal{D}_{tact}$ Fine-tuning
Optimizer type Adam with weight decay Adam with weight decay Adam
Base learning rate 1.5e-4 1.5e-4 1e-4
Weight decay 0.05 0.05 N/A
Learning rate scheduler type Cosine annealing Metric monitoring Cosine annealing
Learning rate reducing factor 0.5 0.5 0.5
Learning rate reducing patience N/A 10 10
First momentum estimate ($\beta_{1}$) 0.9 0.9 N/A
Second momentum estimate ($\beta_{2}$) 0.95 0.95 N/A
Warmup epoch 200 200 10
Total epoch 2000 2000 2000
Early stopping patience 200 200 200
Dropout rate N/A N/A 0.5
Sensitivity loss weight ($\lambda_{s}$) N/A N/A 0.3
We divided the training dataset into 180 groups, each containing 2,500 samples. The three algorithms (PTET, PTET-SL, and SOTA DNN) were trained using subsets of the data consisting of 1, 5, 20, and 180 groups, respectively. The trained models were evaluated on a test dataset comprising 25,000 samples (see test loss in Fig. 4). The PTET model achieved a test loss of 0.002285 using only 2,500 annotated samples. Remarkably, this matches the performance of the SOTA algorithm, which required 450,000 annotated samples (180 times of PTET) to achieve a test loss of 0.002329. Furthermore, PTET significantly outperformed PTET-SL when using less annotated samples (e.g., below 50,000 samples).
Figure 4: Test loss with different numbers of annotated samples.
IV-C Quantitative evaluation
We evaluated the tactile reconstruction performance using three quantitative metrics: Relative Error (RE), Peak Signal-to-Noise Ratio (PSNR) and Correlation Coefficient (CC). These results, summarized in Table III, provide a comprehensive assessment of tactile reconstruction quality from different perspectives. All metrics were calculated as the mean values over the entire testing dataset (25,000 samples), ensuring a robust and reliable evaluation of model performance. Across all metrics, PTET demonstrated superior performance. Remarkably, PTET achieved superior performance using only 0.56% of the annotated samples required by the SOTA DNN model (2,500 vs. 450,000) or just 5% (2,500 vs. 50,000) of the annotated samples required by the PTET-SL, while consistently delivering better results. This highlights PTET’s efficiency and effectiveness in scenarios with limited annotated data, a significant advantage in real-world applications where labelled datasets are scarce or costly to obtain.
To further demonstrate the superiority of PTET, we visually compared the reconstruction results for five representative phantoms from the 25,000 test samples (Fig. 5). The results highlight PTET’s ability to accurately reconstruct positions, shapes and magnitudes of touches across varying touch scenarios, including complex cases with multiple touch points (1–5 touch points).
In contrast, the SOTA DNN, even with 450,000 annotated samples, struggled to capture fine details in scenarios with higher touch complexity, resulting in inaccuracies and noise artefacts. Similarly, PTET-SL trained with 50,000 annotated samples failed to match PTET’s precision and noise reduction capabilities. The self-supervised PTET model, trained with only 2,500 annotated samples, consistently outperformed both alternatives, delivering sharper, more accurate reconstructions and further validating the effectiveness of the self-supervised learning approach.
TABLE III: Quantitative Metrics
Quantitative Metrics RE PSNR CC MSE (×$\times 10^{-2}$)
SOTA (2,500) 0.6479 31.1579 0.6203 0.4049
SOTA (12,500) 0.5811 32.2255 0.6950 0.3220
SOTA (50,000) 0.5213 33.1799 0.7490 0.2663
SOTA (450,000) 0.4941 33.6880 0.7773 0.2329
PTET-SL (2,500) 0.7533 29.6557 0.5128 0.5957
PTET-SL (12,500) 0.6440 31.2573 0.6597 0.4422
PTET-SL (50,000) 0.5309 33.8749 0.7627 0.2981
PTET (2,500) 0.4936 34.3308 0.7985 0.2285
Note: The best results are highlighted in bold.
Figure 5: Representative tactile reconstruction results. For better visualization, all results were refined (Gaussian filtering) and normalized.
V Sensor Fabrication and Characterization
This section describes the EIT-based tactile sensor design, fabrication and characterization, which is utilized to further validate the real-world performance of PTET.
V-A Sensor Fabrication
Figure 6: Sensor fabrication process. (a) Hydrogel fabrication process. (b) Steps of sensor fabrication. (c) Fabrication process of the conductive layer.
We designed a dual-conductivity layer tactile sensor combining a hydrogel layer and a carbon-black/graphite (CBG) conductive layer, leveraging the complementary properties of each material. Hydrogels are widely utilized in electronic skin applications due to their preferable electrical properties, flexibility and biocompatibility [44, 28]. The CBG layer with lower resistance improves EIT-based sensor performance.
The hydrogel precursor synthesis (Fig. 6a) was adapted from a formulation in [44], and optimized for tactile sensing. The hydrogel is primarily composed of acrylamide, a hydrophilic monomer known for polyacrylamide-based hydrogels, which retains water and ensures sensor functionality in humid conditions. The inclusion of lithium chloride, a common ionic conductor, further enhances the hydrogel’s ionic conductivity. The synthesis procedure is as follows:
  1. 1.
    The solution contained 21.5 weight percent (wt%) lithium chloride (Sigma-Aldrich), 8 wtacrylamide (Sigma-Aldrich), 32.27 wtdeionized water and 37 wtethylene glycol (Sigma-Aldrich) was prepared. Ethylene glycol maintains hydrogel pliability across a wide temperature range, essential for stable sensor performance.
  2. 2.
    The solution was stirred at 60°C and 700 rpm for two hours to achieve uniform mixing and dissolution.
  3. 3.
    To initiate a polyacrylamide-based network, 1 wtpolyacrylamide (Sigma-Aldrich) was introduced, providing mechanical stability and enhancing sensor durability.
  4. 4.
    The solution was stirred overnight at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete dissolution and formation of a uniform polymer network.
  5. 5.
    A crosslinking agent, 0.15 wtN,N’-methylenebisacrylamide (Sigma-Aldrich), was added to control the mechanical strength of the gel. Additionally, 0.08 wt2-hydroxy-2-methylpropiophenone (Sigma-Aldrich), a photoinitiator, was added to enable UV curing of the hydrogel.
  6. 6.
    Finally, the mixture was shielded from light and stirred for an additional two hours at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete homogenization.
The CBG conductive layer (Fig. 6c) was fabricated by combining carbon black and graphite. This blend leverages the high surface area and conductivity of carbon black along with the structural integrity and enhanced conductivity of graphite. The fabrication procedure is as follows:
  1. 1.
    3 g of carbon black (Imerys) and 2 g of graphite (Imerys) were dispersed in 23 g of Isopropyl Alcohol (IPA, RS Components) to create a uniform CBG mixture. IPA acts as a dispersing agent, ensuring even distribution of carbon black and graphite particles.
  2. 2.
    Ecoflex 00-30, a silicone elastomer chosen for its superior flexibility, was used as the substrate for the conductive layer. A 1:1 mixture of parts A and B of Ecoflex 00-30 (20 g each, Smooth-On Inc.) was stirred for 3 minutes to initiate cross-linking.
  3. 3.
    20 g of the silicone solution was added to the CBG mixture and stirred for three minutes, followed by degassing to remove any trapped air bubbles.
  4. 4.
    The remaining silicone solution was poured onto a glass plate, and a micrometer-adjustable applicator (VF1823, TQC Sheen) was used to achieve a uniform layer thickness of 0.1 mm. The film was allowed to degas for 10 minutes at room temperature before being cured in an oven at 100 ${}^{\circ}\mathrm{C}$ for 3 minutes.
  5. 5.
    The prepared CBG solution was applied over the cured silicone layer, again using the micrometer-adjustable applicator to maintain a uniform thickness of 0.1 mm. The film was degassed at room temperature for 10 minutes and subsequently cured in an oven at 100 ${}^{\circ}\mathrm{C}$ for another 3 minutes.
  6. 6.
    The cured composite film was cut into a 120 mm $\times$ 120 mm square using a laser cutter.
Finally, the hydrogel and CBG layers were integrated to create the tactile sensor (Fig. 6b). To ensure the sensor’s durability, the overall structure was encapsulated with a silicone layer to prevent moisture loss. The sensor dimensions are 120 $\times$ 120 $\times$ 10 mm3. The assembly procedure is as follows:
  1. 1.
    A 3D-printed mold, containing designated regions for the sensing area, cable channels, and electrode holes, was prepared.
  2. 2.
    Cables and electrodes were installed within the mold.
  3. 3.
    The back of the mold was sealed with a layer of silicone to minimize water loss from the hydrogel due to air exposure.
  4. 4.
    The hydrogel precursor solution was poured into the sensing area and subsequently cured under UV light (365 nm) for two hours to form a stable hydrogel layer.
  5. 5.
    The cured hydrogel layer was then combined with the CBG conductive layer, forming the primary sensing structure.
  6. 6.
    A layer of silicone was applied over the entire sensor assembly to provide environmental sealing and prevent external interference.
The dual-layer design combines the flexibility and biocompatibility of the hydrogel with the enhanced conductivity and structural integrity of the CBG layer, maximizing signal response under applied pressure and making it well-suited for applications in electronic skin and human-machine interfaces.
V-B Sensor Characterization
To evaluate the sensitivity of the developed sensor, we conducted an experiment where the sensor was subject to a continuously increasing localized pressure with strains ranging from 0 mm to 5 mm, as illustrated in Fig. 7. The results demonstrate a clear increase in the conductivity distribution as the compressive strain increases. Conductivity distribution images were generated for compressive strains of 1 mm, 2 mm, and 3 mm. At a compressive strain of 1 mm, the changes in conductivity are subtle yet detectable, while at 2 mm, the effect becomes more pronounced. These results demonstrate the sensor’s preferred sensitivity and capability for accurate and responsive tactile detection.
Figure 7: Sensor sensitivity analysis. The conductivity distribution is generated using Tikhnov Regularization [45].
To evaluate the sensor’s durability and repeatability, we performed a cycling test involving 300 pressing cycles with compressive strains from 2 mm to 5 mm. The results show that the sensor maintains stable output throughout the cycles, with consistent trends in signal changes (Fig. 8). During this test, a dynamic baseline approach was adopted to mitigate slight signal drifts attributed to the lagging effect of the sensor after prolonged pressing. These results confirm the sensor’s robustness and suitability for applications requiring long-term, repeated tactile interactions.
Figure 8: Sensor cycling test for 300 pressing cycles.
VI PTET performance in physical experiments
VI-A Real-world Dataset
A diverse real-world dataset was created using 84 custom-fabricated touch tips produced by a 3D printer. The design of the touch points, including their positions and sizes, was randomly generated, with pressing depths precisely controlled between 2 mm and 5 mm. Data collection was performed using a three-axis linear robot (IGUS, DLE-RG-0001) to ensure precise control over the position and depth of the presses. An in-house developed EIT system was utilized for data acquisition as described in [46].
We collected a total of 5700 annotated samples, distributed as follows: 613 samples with 1 touch point, 1,063 with 2 touch points, 1,043 with 3 touch points, 1,004 with 4 touch points, 995 with 5 touch points, 493 with annular touch and 489 with L-shaped touch. Each sample consists of a 104-dimensional vector of voltage measurements, which is transformed into a 64x64 E2IM matrix and a 48x48 image representing the press map. The samples were randomly shuffled and split into three datasets in a 7:2:1 ratio for training, validation, and testing. To expedite training, we reused the pre-trained models from the simulation phase for both the PTET algorithm and the baseline DNN algorithm. This approach is justified because the pretraining model only captures the fundamental features of voltage measurements and pressure maps.
VI-B Quantification of Reconstruction
We conducted a quantitative evaluation of the test dataset consisting of 570 samples, with results summarized in Table IV. The PTET model demonstrated far superior tactile reconstruction performance across all metrics compared to SOTA DNN. Specifically, PTET reduced the RE by 21.92%, improved the PSNR by 11.15%, increased the CC by 7.96%, and remarkably reduced MSE by 63.88%. To visually compare model performance, we selected seven representative frames from the test Dataset, as displayed in Fig. 9. The DNN model performs comparably to PTET on simpler Phantoms (i.e., Phantoms 1-3), but fails to achieve similar tactile reconstruction quality on more complex Phantoms (Phantoms 4-7). Additionally, we tested model generalization using a set of unseen L-shaped data with sizes not included in the training set. The reconstruction results for this sample (highlighted by red dash-line) achieved quantitative metrics comparable to the test dataset average, demonstrating the strong generalizability of PTET.
Notably, the DNN model chosen for comparison performed exceptionally well in [28], as it was trained on 583,360 annotated simulation samples. However, in our experiments, with only 3,990 annotated data (around 1/150) for training, the DNN model’s performance was significantly inferior. This suggests the superior performance of PTET when trained on a limited amount of annotated data, demonstrating its robustness and efficiency in data-scarce scenarios.
TABLE IV: Average Quantitative Metrics on Test Set
Quantitative Metrics RE PSNR CC MSE (×$\times 10^{-2}$)
PTET 0.3388 32.5171 0.8772 0.1914
SOTA: DNN 0.4339 29.2558 0.8125 0.5299
The best results are highlighted in bold.
Figure 9: Quantification of Reconstruction. The red dash-line shows generalization to an unseen L-shaped sample. Results were refined using Gaussian smoothing.
VI-C Performance on Complex Touches
To evaluate the model’s effectiveness in tactile reconstruction during practical operations, we tested it with a series of complex touch scenarios. These included configurations involving multiple fingers, annular touch, annular touch with finger, and L-shaped touches. Fig. 10 shows the model’s ability to generalize effectively to diverse real-world touch interactions even without specific labels. These results demonstrate the model’s adaptability and robustness for practical applications. Supplementary video 1 shows the real-time tactile reconstruction performance on selected phantoms.
Figure 10: Tactile reconstruction for unseen real-world complex touches. (a) Multi-finger touch, (b) Annular touch, (c) Annular touch with a single finger, and (d) L-shaped touch.
VI-D Comparison of Different Training Datasets
To demonstrate the sim-to-real gap and advantages of using experimental data in tactile reconstruction, we fine-tuned PTET using three datasets: experiment dataset, noise-free simulation dataset and noisy simulation dataset. To ensure fairness, we only consider the circular touch. Each training dataset contains 2500 samples and the validation dataset contains 1000 samples. In addition, we used 504 annotated real-world samples to evaluate performance.
The quantitative metrics (Table V) show that models trained on experimental data significantly outperformed those trained on simulated data or simulated data with noise across all metrics. The experimental data-trained model excelled in reconstructing the number, position, depth, and edge details of touch areas, consistently delivering superior accuracy and detail recovery. While adding noise to the simulated data slightly improved metrics such as CC and PSNR compared to noise-free simulation data, the experimental data still yielded the best overall performance in terms of precision and reconstruction quality.
For visual comparison, Fig. 11 displays reconstructions of five selected phantoms from the test dataset. The results clearly illustrate the model’s superior capability when trained with experimental data, accurately reconstructing the number and spatial distribution of touch points, particularly in depth and edge definition.
TABLE V: Quantitative Metrics of Different Training Datasets
Metrics Experiment Sim (Noise Free) Sim (50dB Noise)
RE 0.3385 0.8589 0.8398
PSNR 34.9600 25.4645 25.6551
CC 0.8633 0.4349 0.4455
MSE (×$\times 10^{-2}$) 0.1551 0.9069 0.8829
Best results are highlighted in bold.
Figure 11: Reconstruction results of different Fine-tuning datasets.
VI-E HMI Applications
To demonstrate real-world applications, the designed tactile sensor is utilized as an HMI interface to control virtual characters in the Google Dino Game and Super Mario Bros Game. Pressing different areas triggers corresponding actions, such as jumping and crouching in Google Dino, while varying press durations enable distinct motion amplitudes, like low- and high-altitude jumps in Super Mario Bros. Additionally, the tactile sensor controls a UR5e robotic arm, allowing it to grab objects at specified locations and arrange them sequentially. These demonstrations highlight the sensor’s ability to capture and transform subtle tactile inputs into real-time, intuitive controls for virtual environments and robotic systems. Supplementary Video 2 showcases real-time interactions with the Google Dino Game, Super Mario Bros Game, and the Universal Robots UR5e.
Figure 12: HMI applications (see supplementary video 2 for real-time control demonstrations).
VII Conclusion
This work introduces the PTET model, a novel framework for EIT-based tactile reconstruction that achieves high-resolution tactile mapping with minimal labelled real-world data through self-supervised pretraining. PTET mitigates the reliance on extensive annotated datasets and requires only 0.56% of the annotated samples compared to the SOTA DNN model (2,500 vs. 450,000) while delivering superior tactile reconstructions. Compared to its supervised counterpart, PTET-SL, PTET achieves comparable performance using just 5% of the annotated samples (2,500 vs. 50,000). Experimental results further demonstrate PTET’s ability to bridge the simulation-to-reality gap and surpass the SOTA DNN model in reconstruction accuracy, achieving up to a 63.88% improvement in overall performance. These findings validate PTET as an effective and scalable solution for tactile perception systems, paving the way for advanced applications in HMI and robotics. Beyond tactile sensing, the pre-trained encoder, based on the EIM representation, can be easily applied to other EIT sensing tasks or serve as a teacher model for developing a versatile EIT foundation model.